OpenStudy (anonymous):

Add. simplify by collecting like radical terms, if possible. Assume that all expressions under radicals represent non-negative numbers 4 sqrt y + 3 sqrt y =

OpenStudy (anonymous):

Okay..so before you can add any radical term you have to make sure that the expression in the radials are the same. In this case they are we have y in both square roots. So all we have to do is to add the coefficient of the square roots which are 4 and 3. \[4\sqrt{y}+3\sqrt{y} = 7\sqrt{y}\]

OpenStudy (anonymous):

got it thanks :-)

OpenStudy (anonymous):

cool!